OpenStudy (anonymous):

Help Medal Indicate which property is illustrated below. 9x + 7y=7y + 9x

OpenStudy (anonymous):

Associative property

OpenStudy (shamil98):

commutative properrty

OpenStudy (shamil98):

commutative property of addition

OpenStudy (anonymous):

thank you @shamil98

OpenStudy (shamil98):

batmun check yo properties bruh http://www.purplemath.com/modules/numbprop.htm

OpenStudy (anonymous):

I guessed

OpenStudy (anonymous):

lolol, but thank you

OpenStudy (shamil98):

that was my guess too lol but i had to check..

OpenStudy (anonymous):

I can never remember the names >_>

OpenStudy (anonymous):

Indicate which property is illustrated in Step 2. Step 1 2x + 11 + 7x + 3 = (2x + 11) + (7x + 3) Step 2 = 2x + (11 + 7x) + 3 Step 3 = 2x + (7x + 11) + 3 Step 4 = (2x + 7x) + (11 + 3) Step 5 = (2 + 7)x + (11 + 3) Step 6 = 9x + 14

OpenStudy (anonymous):

could you all help me out with this one

OpenStudy (shamil98):

that's associative i believe

OpenStudy (anonymous):

which property

OpenStudy (shamil98):

a + (b + c) = (a + b) + c associative property of addition

OpenStudy (anonymous):

Associative